Identify the surfaces of the following equations by converting them into equations in the Cartesian form.


ρ = sin ϕ sin θ


1
Expert's answer
2022-06-06T15:13:51-0400

The equation "\u03c1 = sin \u03d5 sin \u03b8" is in spherical coordinates.

To convert from spherical coordinates to rectangular coordinates there are the next equations:

"x=ρsin\phi cosθ,"

"y=ρsin\phi sinθ,"

"z=ρcos\phi,"

"\u03c1^2=x^2+y^2+z^2."


From the second equation we have: "sin ϕ sin θ=\frac{y}{\rho},"

and we can rewrite the equation of the surface:

"\rho=\frac{y}{\rho},"

"\rho ^2=y,"

"x^2+y^2+z^2=y,"

"x^2+y^2-y+z^2=0,"

"x^2+(y-\frac{1}{2})^2-{\frac{1}{4}}+z^2=0,"

"x^2+(y-\frac{1}{2})^2+z^2= \frac{1}{4}."

This is the equation of a sphere with a center in "(0, \frac{1}{2}, 0)" and radius "\frac{1}{2}."